# Do You Get the Hint? Benchmarking LLMs on the Board Game Concept

### v1

@@ Abstract
Large language models (LLMs) have achieved striking successes on many benchmarks, yet recent studies continue to expose fundamental weaknesses. In particular, tasks that require abstract reasoning remain challenging, often because they use representations such as grids, symbols, or visual patterns that differ from the natural language data LLMs are trained on. In this paper, we introduce Concept, a simple word-guessing board game, as a benchmark for probing abductive reasoning in a representation that is much closer to LLM pre-training data: natural language. Our results show that this game, easily solved by humans (with a success rate of over 90%), is still very challenging for state-of-the-art LLMs (no model exceeds 40% success rate). Specifically, we observe that LLMs struggle with interpreting other players’ strategic intents, and with correcting initial hypotheses given sequential information updates. In addition, we extend the evaluation across multiple languages, and find that the LLM performance drops further in lower-resource languages (Dutch, French, and Spanish) compared to English.
@@ 1 Introduction
While Large Language models have propagated impressive progress in NLP recently, it remains a point of debate how much knowledge they possess about the world, and how (or if at all) they reason about this knowledge (see for example Chollet et al. (2025); Kaesberg et al. (2025); Ngo et al. (2025); Benchekroun et al. (2023); Dentella et al. (2024)). A creative way to assess such capabilities is to have LLMs play (natural language) games. In this paper, we add to this line of research by introducing the boardgame ‘Concept’ as a task to measure LLMs’ abstract and abductive reasoning skills. Similar to a single-modal game of Pictionary, there are two roles in Concept (see Figure 1 for an example):
Player 1, the ‘clue-giver’, chooses a concept (a word or a short phrase) and uses a predefined set of clues to express this concept to Player 2. Player 1 has to come up with creative ways to combine these constrained clues so Player 2 can reconstruct the concept in a meaningful way.
@@
Player 2, the ‘guesser’, attempts to guess the concept based on the clues that are given. In a game, there can be multiple guessers.
Figure 1: A simplified example of a round of Concept, comparing guesses made by humans and an LLM.
@@
Both players have a shared goal, namely for Player 2 to guess the correct concept. Both Player 1 and Player 2 have to strategically navigate the shared conceptual space, created by the constrained vocabulary introduced by the predefined set of clues. Making use of this predefined set, the clue-giver has the freedom to express the concept however they see fit. Therefore, the process of clue-giving is inherently more imprecise, complicating a fair evaluation of LLMs (i.e., does the model not know the concept, or is it unable to select and combine seemingly unrelated clues in a meaningful way?). Therefore, in our setup, we implement LLMs solely as the guesser, using human-generated clues as input.
@@
    Are some categories of concepts (e.g., animals, objects, quotes) easier for models?
  3. 3.
@@
    We expect that model performance does not match, but correlates with human performance. Additionally, we expect that reasoning models outperform non-reasoning models on this task.
@@
    LLMs perform better on everyday and concrete concepts (e.g., objects, animals) compared to infrequent or abstract concepts (e.g., quotes, culturally-dependent concepts).
  3. 3.
    Performance on low-resource languages is worse.
Our results show that, similar to observations made on visual and spatial reasoning tasks such as ARC and SPaRC, LLMs struggle on a simple language game that humans solve easily (94% solve rate for non-expert humans). Specifically, across all tested models, we observe difficulties with (i) interpreting other players’ strategic intent from the clues they provide, and (ii) correcting initial hypotheses when new information becomes available.
@@ 2 Related Work
The core challenge that is addressed in this work relates to logical inference and abductive reasoning. Historically, this task was approached by creating symbolic knowledge representations for AI, but has ultimately failed to generalize because of the brittleness and incompleteness of knowledge bases. Later, with the rise of representation learning, distributed vector embeddings became the dominant method for encoding knowledge, enabling models to capture semantic relationships in large text corpora (Mikolov et al., 2013). Building on this foundation, the rise of large autoregressive pre-trained transformer models has made it possible to leverage natural language itself as the medium of knowledge representation (Yang et al., 2023). This evolution has opened the way for creative probes to test reasoning skills in LLMs.
@@
Game settings offer interesting perspectives to test AI models’ capabilities, and understanding of a (physical) world. The rise of powerful auto-regressive Large Language Models (LLMs) gave a significant boost to this research direction. LLMs can embody various roles in games: they can be players, non-player characters, assistants, Game Masters, design games, or analyze gameplay data, and so on (Gallotta et al., 2024). (Natural language) games have been used to evaluate core skills of LLMs, in addition to the standardized benchmarking practices in NLP. For instance, situated language understanding is evaluated through dialogue games (Chalamalasetti et al., 2023), theory of mind is tested through collaborative games (Li et al., 2023a), physical common sense is probed by asking models to organize rooms and place objects in appropriate locations (Zhuo and Murata, 2024) and world knowledge can be tested through the popular game ‘Twenty Questions‘ (De Bruyn et al., 2022). It has also been shown that LLMs as players do not necessarily match human performance, but their performance correlates strongly with difficulty as indicated by humans (Xiao and Yang, 2024).
@@
An interesting subsection of games that are implemented for LLMs are games in which there is a perfect common interest between players. In this setup, there are usually two players: the ‘sender’, and ‘receiver’. The sender disposes of a fixed set of signals. The receiver can observe the signals that are sent, but not the state of the world. A message is encoded in signals by the sender, and the receiver must then try to decode them and decide on which actions to take. For each state, there is only one correct action and both the sender and the receiver have the same outcome objective, namely for the receiver to perform the correct action. Prior research has presented such games as a method to evaluate automated methods, since they test a system’s ability to strategically encode and/or decode information, align meaning between agents, and coordinate actions within a shared framework.
@@
For instance, LLMs have made considerable gains in solving crossword puzzles, especially with the implementation of search algorithms (Saha et al., 2024). However, cryptic puzzles (e.g., relying on wordplay) are much more challenging, and there remains a large performance gap compared to humans (Saha et al., 2024; Sadallah et al., 2024). Leng et al. (2025) create a text+grid crossword benchmark, and show that reasoning models outperform non-reasoning models. However, they do not mention a human baseline. Similarly, the game ‘Codenames’ has been introduced as benchmark to test language nuances and theory of mind, on which LLMs again outperform older methods. However, Stephenson et al. (2025) note that smaller models do not adhere to the constraints of game rules, and models struggle overall with lateral thinking. On the popular puzzle ‘Wordle’, LLMs again underperform but CoT-prompting increases performance (Xiao and Yang, 2024).
@@
However, these games are based on static situations. Lately, interactive games have been used to gauge models’ continuous understanding updates, context-aware responses, and the inference of hidden states of other agents (Momentè et al., 2025). It is argued that these games offer a better proxy for the real-world necessity of adaptability and robustness in interactions. Such game-based evaluations have demonstrated higher correlations with a range of cognitive abilities, including causal and logical inference, as well as social and emotional competencies (Momentè et al., 2025). For instance, in ‘Guess The Build’, players construct a Minecraft build and others try to guess it with natural language. Results show a considerable performance gap between open and API Visual Language Models (VLMs). The performance correlates with frequency of the concept in the training data, and underrepresented cultural concepts and low-resource languages remain an open problem (Zhu et al., 2025). Similarly, the game ‘Pictionary’ has been adapted as a multi-modal benchmark ‘Iconary’, in which a guesser tries to identify a concept that is drawn by composing a set of icons, and the drawer can update their drawing to help the guesser. Results show that T5-models are better at guessing compared to drawing, and the main challenge were out-of-vocabulary words (Clark et al., 2021). Pictionary has also been played by VLMs using the dataset and agentic framework ‘Sketchtopia’, where the win rate is lower compared to humans, and the performance is again linked to frequency in training data (Khan and Sarvadevabhatla, 2025).
@@
In order to be successful in such games, being able to estimate mental states of other players (such as their knowledge, beliefs, or desires), known as Theory of Mind (ToM), is important. While Kosinski (2024) observes that these skills are improving with newer and larger LLMs, Akata et al. (2025) find that LLMs often fail in game settings where cooperation with other players is required, suggesting that ToM understanding of other players is limited. Similarly, Li et al. (2023b) find long contexts and hallucinations are barriers to ToM in LLMs for multi-agent collaboration in text games. Agashe et al. (2023) show that LLMs "perform well in multi-agent coordination settings where decision-making primarily relies on environmental variables but face challenges in scenarios requiring active consideration of partners’ beliefs and intentions".
We contribute to this line of research by introducing ‘Concept’ as a benchmark to evaluate abductive reasoning in LLMs. In this game, the sender selects clues to maximize the probability of the receiver correctly identifying the concept, while constrained to the set of available clues. The receiver, in turn, strategically interprets these symbolic clues, inferring the most plausible concept that best satisfies all of them. The effectiveness of the clues relies on how well the sender’s intended meaning aligns with the receiver’s interpretation. The players must converge on a shared understanding of the concept based on the limited, symbolic clues. This highlights the importance of not just linguistic comprehension but also the strategic intent behind clue selection and interpretation. In order to perform well in Concept, it is not enough for AI to understand each clue separately, but it must be able to meaningfully combine these clues, within the shared conceptual framework that humans use.
‘Concept’ has the advantage of being a very straightforward task, for which there is no need for expert knowledge. Additionally, our setup allows us to compare performance with a human baseline, and the inclusion of multiple languages offers insights into the transferability of LLM’s abduction capabilities. In this way, our benchmark complements popular reasoning challenges such as ARC (Chollet et al., 2025) and SPaRC (Kaesberg et al., 2025). The ARC (Abstraction and Reasoning Corpus) benchmark evaluates generalization through visual pattern-recognition puzzles that require discovering abstract rules from small sets of examples, while SPaRC (Spatial Pathfinding Reasoning Challenge) builds on this idea with text-based grid environments that demand spatial reasoning and step-by-step planning to navigate toward a goal. Both benchmarks probe general reasoning skills but rely on visual or spatial representations that differ substantially from the linguistic data on which LLMs are trained. In contrast, because Concept relies solely on natural language, its representation is more closely aligned with the pre-training data of LLMs, thereby providing a fairer test of their reasoning capabilities.
@@ 3.1 Data
We collected game logs of Concept from the online boardgame platform Board Game Arena (BGA) [^3] https://en.boardgamearena.com/gamepanel?game=concept. Only completed games were kept, and all player data was anonymized (i.e., we removed player ID, timing, stats, etc.). For each language in our dataset (English, French, Dutch, and Spanish), we collected 100 game logs, which amounts to circa 1,000 concepts played per language. In this section, we will detail the structure of our dataset.
Each game includes multiple rounds, with one concept per round. On average, one game consists of 12 concepts. While each concept within a single game is unique, the same concept can appear across different games; in our dataset, the most frequent concept occurs nine times within one language. Importantly, even when the target concept is repeated, the clues differ, since they are provided by different players in different games.
@@
On the collected dataset, we filter out the rounds for which there were no clues given (i.e., Player 2 guessed the concept before Player 1 gave any clues). In addition, some rounds contain very few clues because Player 1 removed most of them at the end of the game. To account for this, we report results on two versions of the dataset: one containing all rounds, and one excluding incomplete rounds. Our filtering heuristic considers a round incomplete if the initial clue (the top-level and most important clue) is missing at the end of the game. Table 1 provides an overview of the resulting dataset sizes.
@@ 3.2 Prompting
We include a wide range of LLMs, both open and closed-sourced: LlaMA3.3 70B; LlaMA4-Maverick; Qwen2.5 72B; Deepseek V3; Mistral Small 24B; and GPT 4.1 mini. Additionally, we include the GPT-OSS 120B reasoning model. These models were chosen because they are among the most recent high-performing releases across major model families, spanning diverse architectures and parameter scales, offering a representative overview of current LLM capabilities. In Appendix A, we include more model details.
@@
For non-reasoning models, we allow 10 output tokens, and keep the temperature stable at 0.1 to keep the output more deterministic and reproducible. For GPT-OSS-120B, we allow 1,000 output tokens, with medium reasoning effort, and follow the recommended temperature setting (t=1). Because of hardware limitations, we prompt the models using the Together.ai API [^4] https://www.together.ai/, except for GPT 4.1 mini, for which we use the OpenAI API [^5] https://platform.openai.com/docs/models/gpt-4.1-mini. The total cost for prompting the models was 120\$. We keep the prompt identical for all models. We organize the clues from Player 1 in a hierarchical structure (grouping clues of the same color together), since tests on the separate validation set (n=50) showed that this technique results in better performance. There is no limit to the number of clues that can be given, as is the case in the original human game. We also include one example in the prompt, taken from the rule book of the original game. Prior works found that one example can work better than multiple examples, ‘possibly due to cognitive overload or excessive focus on example analysis instead of task solving’ (Kaesberg et al., 2025; Ye et al., 2023).

### v2

@@ Abstract
Large language models (LLMs) have achieved striking successes on many benchmarks, yet recent studies continue to expose fundamental weaknesses. In this paper, we introduce Concept, a simple word-guessing board game, as a benchmark for probing abductive reasoning. Our results show that this game, easily solved by humans (with a success rate of over 90%), is still very challenging for state-of-the-art LLMs (no model exceeds 40% success rate). Specifically, we observe that LLMs struggle with interpreting other players’ strategic intents, and with correcting initial hypotheses given sequential information updates. In addition, we extend the evaluation across multiple languages, and find that the LLM performance drops further in lower-resource languages (Dutch, French, and Spanish) compared to English.
@@ 1 Introduction
While Large Language models have propagated impressive progress in NLP recently, it remains a point of debate how much knowledge they possess about the world, and how (or if at all) they reason about this knowledge (e.g. chollet2025arc; kaesberg2025sparc; ngo2025mscore; benchekroun2023worldsense; dentella2024testing; gevers-etal-2025-benchmarks). A creative way to assess such capabilities is to have LLMs play (natural language) games. In this paper, we introduce the boardgame Concept as a task to measure LLMs’ abstract and abductive reasoning skills. We illustrate the construction of our benchmark in Figure 1. Similar to a single-modal game of Pictionary, there are two roles in Concept:
Figure 1: The core components of Concept. Step 1: we collect human game logs. Step 2: we extract the clues from Player 1 to create the prompts, which are then given to the LLM. In Step 3 we compare the performance between humans and LLMs on the task.
Player 1, the ‘clue-giver’, chooses a concept (a word or a short phrase) and selects clues from a predefined list to express this concept to Player 2. Clues are added by using a pawn or cube of a specific color: clues of the same color are (hierarchically) related to each other. The first clue establishes the main category of the concept and is marked by a green question mark pawn (see Figure 1). Player 1 can give clusters of clues that relate to this main category by introducing exclamation mark pawns of different colors. The low-level clues (cubes) help to clarify the high-level clue (pawn) of the same color. Thus, low-level clues should be interpreted in relation to the high-level clue of that same color. For instance, in the example in Figure 1, the green question mark indicates the highest-level category (‘Building’), where the small green hint (’Metal’) refers to the material of the building. The red exclamation mark marks a high-level clue describing that building (‘Flag’), and the low-level red clues (‘Blue’, ‘White’, and ‘Red’) should be understood as properties of that flag, not of the building directly. In this case, the clues mean ‘a metal building located in France’.
@@
Player 2, the ‘guesser’, attempts to guess the concept based on the clues that are given. In a game, there can be multiple guessers. The only feedback Player 2 gets, is whether their answer is correct or not. Player 1 can in turn update their clues based on the incorrect answers, until Player 2 finds the correct answer.
We introduce Concept as a benchmark to evaluate abductive reasoning in LLMs. In this game, Player 1 selects clues to maximize the probability of the receiver correctly identifying the concept, while constrained to the set of available clues (118 unique items) [^1] The icons with their textual representation can be found in the rule book, e.g., https://shorturl.at/pWW0D. Player 2, in turn, has to use the available (incomplete) and abstract evidence (i.e., the clues) to efficiently search over a very open semantic space, to infer the most plausible concept that best satisfies all of the clues. To do so, Player 2 must rely on their own implicit background knowledge, which helps them to correctly interpret the clues (for instance, in the example in Figure 1, knowing that the clue ‘Metal’ in relation to ‘Building’ refers to the material the building is made of). Since multiple answers could theoretically be correct given the clues, the player must also take into account the strategic intent of the clue-giver (relating to their Theory of Mind), and learn from previous incorrect answers.
@@
Because the clue-giver has the freedom to express the concept however they want (while constrained to the available clues), this process of clue-giving is inherently more imprecise. This complicates a fair evaluation of LLMs (i.e., does the model not know the concept, or is it unable to select and combine seemingly unrelated clues in a meaningful way?). Therefore, in our setup, we implement LLMs solely as the guesser, using human-generated clues as input.
The effectiveness of the clues relies on how well the sender’s intended meaning aligns with the receiver’s interpretation. The players must converge on a shared understanding of the concept based on the limited, symbolic clues. This highlights the importance of not just linguistic comprehension, but also the strategic intent behind clue selection and interpretation. In order to perform well in Concept, it is not enough for the model to understand each clue separately, but it must be able to meaningfully combine these clues, within the shared conceptual framework that humans use.
Evaluating an LLM’s performance on Concept thus assesses its capacity to participate such strategic linguistic interaction, relying on abductive reasoning that integrates associative skills (i.e., to disambiguate and connect the clues), Theory of Mind (i.e., to infer the strategic intention of Player 1), and common sense (i.e., to supply the (often unstated) background facts and everyday associations needed to correctly interpret the clues).
Concept has the advantage of being a very straightforward task, for which no expert knowledge is needed. Additionally, our setup allows us to compare performance with a human baseline, and the inclusion of multiple languages offers insights into the transferability of an LLM’s abduction capabilities.
@@
    We expect that model performance does not match, but correlates with human performance. Additionally, we expect that reasoning models outperform non-reasoning models, and that everyday and concrete concepts (e.g., objects, animals) are easier for models compared to infrequent and abstract concepts (e.g., quotes).
@@
    Performance on low-resource languages is worse than on English.
Our results show that LLMs struggle on a simple language game that humans solve easily (94% solve rate for non-expert humans). Specifically, across all tested models, we observe difficulties with (i) interpreting other players’ strategic intent from the clues they provide, and (ii) correcting initial hypotheses when new information becomes available.
@@ 2 Related Work
As mentioned earlier, the core challenge that is addressed in this work relates to logical inference and abductive reasoning. The guesser must interpret the (incomplete) abstract clues to infer the most plausible concept, relying on their own background knowledge, Theory of Mind of the other player, and previous incorrect guesses. Language games offer an interesting perspective to test these capabilities in AI models. For instance, situated language understanding is evaluated through dialogue games (chalamalasetti2023clembench), Theory of Mind (ToM) -the capacity to reason about others’ concealed mental states- is tested through collaborative games (li-etal-2023-theory), physical common sense is probed by asking models to organize rooms and place objects in appropriate locations (zhuo-murata-2024-utilizing), and world knowledge can be tested through the popular game ‘Twenty Questions‘ (de-bruyn-etal-2022-20q). It has also been shown that LLMs as players do not necessarily match human performance, but their performance correlates strongly with difficulty as indicated by humans (xiao2024llms).
@@
An interesting subsection of games that are implemented for LLMs are games in which there is a perfect common interest between players. In this setup, there are usually two players: the ‘sender’, and ‘receiver’. The receiver can observe the signals that are sent, but not the state of the world. A message is encoded in signals by the sender, and the receiver must then try to decode them and decide on which actions to take. For each state, there is only one correct action and both the sender and the receiver have the same outcome objective, namely for the receiver to perform the correct action. Prior research has presented such games as a method to evaluate automated methods, since they test a system’s ability to strategically encode and/or decode information, align meaning between agents, and coordinate actions within a shared framework.
@@
For instance, LLMs have made considerable gains in solving crossword puzzles, especially with the implementation of search algorithms (saha2024language). However, cryptic puzzles (e.g., relying on wordplay) are much more challenging, and there remains a large performance gap compared to humans (saha2024language; sadallah2024llms). leng2025crosswordbench create a text+grid crossword benchmark, and show that reasoning models outperform non-reasoning models. However, they do not mention a human baseline. Similarly, the game ‘Codenames’ has been introduced as benchmark to test language nuances and ToM, on which LLMs again outperform older methods. However, stephenson2025codenames note that smaller models do not adhere to the constraints of game rules, and models struggle overall with lateral thinking. On the popular puzzle ‘Wordle’, LLMs again underperform but CoT-prompting increases performance (xiao2024llms). Additionally, the ‘Connections’ puzzle in the New York Times, where players have to make four clusters from sixteen words, is still challenging for LLMs. Especially when the model’s first answer is incorrect, the accuracy is low (todd2024missedconnectionslateralthinking). Similarly, in Italian, the game ‘La Ghigliottina’ in which a single player has to determine which concept connects a set of five seemingly unrelated clues, shows that LLMs are still underperforming (manna-etal-2024-riddle).
@@
However, these games are based on static situations. Lately, interactive games have been used to gauge models’ continuous understanding updates, context-aware responses, and the inference of hidden states of other agents (momente2025triangulating). These dynamic, or interactive games provide an information exchange between the sender and receiver. It is argued that these games offer a better proxy for the real-world necessity of adaptability and robustness in interactions. Such game-based evaluations have demonstrated higher correlations with a range of cognitive abilities, including causal and logical inference, as well as social and emotional competencies (momente2025triangulating). For instance, in ‘Guess The Build’, players construct a Minecraft build and others try to guess it in natural language. Results show a considerable performance gap between open and API Visual Language Models (VLMs). The performance correlates with frequency of the concept in the training data, and underrepresented cultural concepts and low-resource languages remain an open problem (zhu2025guessbenchsensemakingmultimodalcreativity). Similarly, the game ‘Pictionary’ has been adapted as a multi-modal benchmark ‘Iconary’, in which a guesser tries to identify a concept that is drawn by composing a set of icons, and the drawer can update their drawing to help the guesser. Results show that T5-models are better at guessing than drawing, and the main challenge were out-of-vocabulary words (clark-etal-2021-iconary). Pictionary has also been played by VLMs using the dataset and agentic framework ‘Sketchtopia’, where the win rate is lower compared to humans, and the performance is again linked to frequency in training data (khan2025sketchtopia).
@@
In order to be successful in such games, being able to estimate mental states of other players (such as their knowledge, beliefs, or desires), i.e., ToM, is important. While kosinski2024evaluating observes that these skills are improving with newer and larger LLMs, akata2025playing find that LLMs often fail in game settings where cooperation with other players is required, suggesting that ToM of other players is limited. Similarly, li2023theory find long contexts and hallucinations are barriers to ToM in LLMs for multi-agent collaboration in text games. Additionally, agashe2023llm show that LLMs "perform well in multi-agent coordination settings where decision-making primarily relies on environmental variables but face challenges in scenarios requiring active consideration of partners’ beliefs and intentions".
Concept adds to this line of research by presenting the task both as a static challenge and a dynamic one, bridging the gap between the two usually separate methodologies. In the static setting, we present the final set of clues, whereas in the dynamic setting we update the clues iteratively. Additionally, our setup allows us to compare LLM performance directly to a human baseline, which is often not possible for other games. Moreover, whereas the majority of benchmarks and games are focused on English, our benchmark covers several languages, allowing us to assess abductive reasoning in a genuinely multilingual setting. The core challenge in Concept, inferring the most plausible answer given ambiguous/ incomplete set of hints, has been tackled before in a multimodal approach (e.g., ‘Guess The Build’ and ‘Pictionary’), but to the best of our knowledge not in a single-modal and multilingual setting. This makes the task more straightforward, and aligns the task representation more with the LLMs’ pre-training data.
@@ 3.1 Data
We collected game logs of Concept from the online boardgame platform Board Game Arena (BGA) [^4] https://en.boardgamearena.com/gamepanel?game=concept. We will release the data for research purposes only.. Only completed games were kept, and all player data was anonymized (i.e., we removed player ID, timing, stats, etc.). For each language in our dataset (English, French, Dutch, and Spanish), we collected 100 game logs, which amounts to circa 1,000 concepts played per language. In this section, we will detail the structure of our dataset.
Each game includes multiple rounds, with one concept per round. On average, a single game consists of 12 concepts. While each concept within a game is unique, the same concept can appear across different games; in our dataset, the most frequent concept occurs nine times within one language. Importantly, even when the target concept is repeated, the clues differ, since they are provided by different players in different games.
@@
On the collected dataset, we filter out the rounds for which no clues were given (i.e., Player 2 guessed the concept before Player 1 gave any clues). In addition, some rounds contain very few clues because Player 1 removed most of them at the end of the game. To account for this, we report results on two versions of the dataset: one containing all rounds, and one excluding incomplete rounds. Our filtering heuristic considers a round incomplete if the initial clue (the top-level and most important clue) is missing at the end of the game. Table 1 provides an overview of the resulting dataset sizes.
@@ 3.2 Prompting
We include a wide range of LLMs, both open- and closed-sourced: LlaMA3.3 70B; LlaMA4-Maverick; Qwen2.5 72B; Deepseek V3; Mistral Small 24B; and GPT 4.1 mini. Additionally, we include the GPT-OSS-120B reasoning model. These models were chosen because they are among the most recent high-performing releases across major model families, spanning diverse architectures and parameter scales, offering a representative overview of current LLM capabilities. In Appendix A, we include more model details and hyperparameter settings.
@@
We keep the prompt identical for all models. We organize the clues from Player 1 in a hierarchical structure (grouping clues of the same color together), since tests on the separate validation set (n=50) showed that this technique results in better performance. There is no limit to the number of clues that can be given, as is the case in the original human game. We also include one example in the prompt, taken from the rule book of the original game. Prior works found that one example can work better than multiple examples, ‘possibly due to cognitive overload or excessive focus on example analysis instead of task solving’ (kaesberg2025sparc; ye2023comprehensivecapabilityanalysisgpt3).
We implement two prompts for the main experiments (in English). Specifically, we create a static and a dynamic prompt. As addressed in Section 2, the state of the game can either be given statically, where Player 2 gets all information at once, or dynamically, where the information is updated sequentially and there is an interaction between the two players.
The static prompt consists of the final set of clues given by Player 1 (similar to a snapshot of the clues as they were at the end of the round), and any incorrect guesses made by the LLM are added to the prompt. We allow 10 guesses per round. This informs us on how well LLMs can navigate the conceptual space, created by humans, and how well they can combine symbolic clues to infer the correct answer. We formalize the procedure in pseudocode 1:
Input: $C$ – fixed set of clues
$W\leftarrow[\ ]$;
$\triangleright$ incorrect guesses
1
2while true do
3 $answer\leftarrow\text{GENERATE\_ANSWER}(C+W)$;
4 if IS_CORRECT(answer)(answer) then
5    return answeranswer;
6
7 $W.\text{append}(answer)$;
8
Algorithm 1 Static prompt